OpenStudy (broner33):

Any line with an undefined slope is parallel to the a.x-axis b.y-axis c.line y = x

Nnesha (nnesha):

\[\huge\rm \frac{ y_2 -y_1}{ x_2 -x_1 }=m\] formula to find slope when will you get undefined slope ??

Nnesha (nnesha):

when you have same x or same y values ?? x= number ( vertical line) y= number (horizontal line )

OpenStudy (broner33):

so y=x

Nnesha (nnesha):

y=mx+b is same as y=mx+0 where m is slope and b represent y-intercept

OpenStudy (broner33):

im sorry, i dont get it i suck at this:(

Nnesha (nnesha):

when would be the fraction undefined ?? if you have 0/a or a/0 ?

OpenStudy (broner33):

o/a i think

Nnesha (nnesha):

well you can divide 0 by a number but you can't divide anything by 0 can you >??

Nnesha (nnesha):

\[\large\rm \frac{ anything }{ 0 }=undefined \]

Nnesha (nnesha):

\[\huge\rm \frac{ y_2 -y_1}{ x_2 -x_1 }=m\] formula to find slope this fraction representt the slope formula in order to get undefined slope you need same x values so when you subtract them you will get 0 like 2-2 , 3-3

OpenStudy (broner33):

so it was c?

Nnesha (nnesha):

now draw a line on the graph http://prnt.sc/aa3eue y values are constant when you draw horizontal line and the x values would be constant when we draw vertical lines (that will give you undefiend slope )

Nnesha (nnesha):

y=x is same as y=1x which is same as y=mx where m is the slope

OpenStudy (broner33):

oooo yhanks i get it know its y- axis right?